# Extended Fractional Chern Insulators Near Half Flux in Twisted Bilayer Graphene Above the Magic Angle

### v2

@@ Abstract
Fractional Chern insulators (FCIs)—the lattice analog of the fractional quantum Hall states—form as fractionalized quasiparticles emerge in a partially-filled Chern band. This fractionalization is driven by an interplay of electronic interaction and quantum geometry of the underlying wavefunctions. Bilayer graphene with an interlayer twist near the magic angle of 1.1°  hosts diverse correlated electronic states at zero magnetic field. When the twist angle exceeds 1.3°, the electronic bandwidth is sufficient to suppress the zero-field correlated states. Yet applying a magnetic field can restore the importance of electron-electron interactions. Here, we report strongly-correlated phases when a 1.37° twisted bilayer graphene sample is tuned to near half a magnetic flux quantum per moiré cell, deep into the Hofstadter regime. Most notably, well-quantized odd-denominator FCI states appear in multiple Hofstadter subbands, over unusually large ranges of density. This suggests a mechanism beyond disorder is stabilizing the fractional states. We also observe a bending and resetting of the Landau minifan reminiscent of the cascade of Dirac resets observed in magic-angle samples near integer filling at zero field.
@@ Introduction
Twisted bilayer graphene (TBG), the archetypal strongly-correlated moiré heterostructure, displays an astonishingly diverse set of correlated electronic and topological phases [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. The electronic moiré miniband structure depends sensitively and predictably on the twist angle. Near the magic angle of 1.1°, the large moiré length scale enables tuning the density of charge carriers through entire minibands and the density of magnetic flux to on order one flux quantum $\Phi_{0}=e/h$ per moiré cell, where $e$ is the elementary charge and $h$ is Planck’s constant. In this “Hofstadter’s butterfly” regime [13, 14, 15, 16, 17, 18, 19, 11, 20, 21, 22, 23, 24], the energy spectrum exhibits a fractal structure with field-induced topological subbands. Gaps in the fractal spectrum appear at fluxes and densities described by Diophantine equations of the form $n/n_{s}=s+t\Phi/\Phi_{0}$ [25]. Here, $n_{s}=1/A$, $\Phi=BA$, $A$ is the moiré unit cell area, $B$ is the magnetic field normal to the plane of the samples, $s$ is the density offset at zero flux, and $t$ is the Chern number associated with the gap. We notate these Středa lines as $(s,\ t)$. Within such a gap, the Hall conductance is expected to be quantized to $\sigma_{xy}=te^{2}/h$ [26].
Integer $t$ Středa lines have been observed in TBG at a range of twist angles. Some of these states are not describable by single-particle models, as symmetries are broken by electronic interactions [27, 7, 12, 22, 24]. Středa lines with fractional $t$ and nonzero $s$ are more exotic and are referred to as fractional Chern insulators (FCIs). Analogous to fractional quantum Hall (FQH) states, where $s=0$, FCIs host fractionalized excitations [28]. Only a few examples of such states in moiré materials have been experimentally reported: twisted MoTe2  [29, 30, 31] and multilayer twisted graphene aligned with hexagonal boron nitride (hBN)  [19, 8, 32, 33].
@@
In Refs. [36] and [37], we presented magnetotransport measurements and single-particle calculations for a TBG sample twisted to near 1.37°, where the miniband width is large enough (near 100 meV) to suppress correlated electronic states near zero magnetic field [1, 3, 36]. In this work, we further explore that same sample, now at extremely high magnetic fields, demonstrating many $B$-induced strongly correlated states. Most notably, at magnetic flux near half a magnetic flux quantum per moiré unit cell we observe plateaus in Hall resistance corresponding to ±$\pm 8/3$ and $-8/5$, quantized to within a few tenths of a percent, along with possible $-4/3$ plateaus. The ±$\pm 8/3$ plateaus in particular extend over a strikingly broad range of density, up to half an electron per moiré unit cell, rather than following a narrow, well-defined Středa line. Integer quantized states have been known to extend over a broad range of densities, outcompeting fractional states that might be expected at those densities. This “re-entrant quantum Hall effect” has been explained as a consequence of a Wigner crystal coexisting with a filled Landau level [38, 39], a scenario originally termed a partial Hall crystal [40]. However, in the sample we study here the integer states are less robust and broad than the $8/3$ fractional states. In Sec. V we discuss several mechanisms that might be responsible for the novel and surprising extended FCI regions.
@@ I Longitudinal transport
In the sample we study, longitudinal resistivity as a function of carrier density and magnetic field between 14 and 28 T ranges from below our measurement floor of a few ohms within quantum Hall gaps to hundreds of kilohms at band edges and near charge neutrality (Fig. 1A). Here, we focus on specific contact pairs within this 20-terminal Hall bar that demonstrated the sharpest low-field transport in Ref. [36]. For our twist angle of 1.37°, $\Phi/\Phi_{0}=1$ occurs at nearly 45 T, so the field range in our measurement corresponds to flux ratios between 0.31 and 0.62. We indicate the most prominent Středa lines, $(s,\ t)=(0,\,\pm 4)$, which persist without closing down to near-zero field (Fig. S4).
Figure 2: Fractional Chern insulating Hall plateaus. (A) Longitudinal resistance from Fig. 1 replotted in grey-scale. Superimposed on the longitudinal resistance are locations where the Hall resistance for a neighboring contact pair is quantized to within 1% of ±$\pm 8/3$ in red and blue, respectively. A clustering algorithm has been applied to remove spurious points (see Sec. S3). The abrupt loss of quantization at $\Phi/\Phi_{0}$ between $1/3$ and $2/5$ corresponds to a pause in the measurement of several hours. This, along with the noise above $1/2$, are described in Sec. S2. (B, C) Hall resistance and longitudinal resistivity as a function of density at 18.5 T ($\Phi/\Phi_{0}=0.41$, black dashed lines in A).
For comparison, we show the computed Wannier plot in the same field range (Fig. 1B). The computed density of states $dn/d\mu$ should show patterns similar to those seen in resistivity, but we do not directly calculate transport from our model as we did in [37] for lower magnetic fields. Our model broadly captures the splitting and bending behavior of the Landau levels emitting from half flux and $n/n_{s}=0$, where open orbits from uniaxial heterostrain and Zeeman splitting play the role of anisotropic hopping and energetic splitting as described in Ref. [36]. The model predicts a number of prominent gaps that we discuss in Sec. III.
In the supplementary materials we describe other striking phenomena such as Brown-Zak oscillations, symmetry-broken Chern insulating Středa lines with $s=\pm 1/2$, $t=\pm 3$, and correlated Hofstadter ferromagnets.
@@ II Fractional Chern Insulators
In Fig. 2A, we overlay in blue and red the regions where the Hall resistance is quantized to within 1% of $-3h/8e^{2}$ and $+3h/8e^{2}$ respectively. We apply an intentionally-conservative clustering threshold algorithm so that almost all of the pockets marked with color represent well-quantized plateaus of ±$\pm 8/3$ rather than regions where the Hall resistance incidentally passes through quantized values. These fractional Hall plateaus coincide with low longitudinal resistance (shown in gray). See Sec. S3 for details and Fig. S21 for unfiltered data.
@@
Fig. 2 panels B and C show line cuts of $R_{yx}$ and $R_{xx}$ at 18.5 T. In addition to the aforementioned ±$\pm 8/3$ plateaus, a number of integer plateaus are visible, along with a small $-8/5$ plateau. Strikingly, the ±$\pm 8/3$ plateaus are quantized over larger ranges of density and field than any integer state. For a discussion of the degree of quantization, other fractions, and the behavior in other contact pairs, see Secs. S10,  S11, and  S7, respectively.
@@
Figure 3: Computed Hofstadter spectrum and candidate quantum spin Hall state. (A) Computed energy levels for valley K, $\theta=1.35^{\circ}$ with $0.24\%$ uniaxial heterostrain at 45° and $0.3\%$ biaxial heterostrain ($q\leq 72$), as described in the text. Spin up electrons are shown in blue, and spin down electrons are shown in red. The gaps at $(s,t)=(0,-4)$ and $(4,-4)$ are labeled. $\alpha$, $\beta$, and $\gamma$ indicate the Hofstadter subbands which host the largest pockets of ±$\pm 8/3$ quantization shown in Fig. 2. Because of the Zeeman spin-splitting, it is possible for the net spin of all filled bands associated with a given gap to be non-zero. A trivial gap (t=0) with a net spin is reminiscent of quantum spin Hall. Two such gaps are indicated: (0, ±$\pm 2$). See Fig. S20 for the full range of fields. (B-C) Dependence of the longitudinal resistance (top panels) and Hall resistance (bottom panels) as a function of density and field centered on the two quantum spin Hall candidate gaps predicted from (A). In (B), below half flux, we observe a strong vertical suppression of longitudinal resistance coincident with a small plateau of zero Hall resistance at $n/n_{s}=-2$. In (C), we observe low longitudinal resistance along a vertical line, however the Hall resistance is not zero at the same densities.
@@
Regions where the Hall resistance is (fractionally) quantized typically coincide with suppressed longitudinal resistance in the four adjacent contact pairs (see Fig. 1 and Sec. S7). Though measurements from neighboring longitudinal probes vary subtly, neighboring Hall probes exhibit strikingly different behavior: those just 3 microns away from the main Hall pair we focus on in this manuscript exhibit poor quantization, even for integer quantum Hall states (Fig. S7), possibly due to mixing of the longitudinal resistance into the Hall measurement because of spatial variation in twist angle  [^1] The standard approach to remove the effect of mixing is to (anti)symmetrize the (Hall) longitudinal resistance as a function of magnetic field. Antisymmetrization could help reveal well-quantized plateaus in other Hall pairs. Given the time constraints of our measurement run at the National High Magnetic Field Lab, we were unable to acquire data at negative fields.. We therefore suspect that the main Hall pair of this manuscript contacts a region of unusually high moiré uniformity.
@@ III Hofstadter Model
To understand the Hofstadter bands we would expect without interactions we perform bandstructure computations within a single-particle model. Our effective continuum model from Ref. [37] which accounted for uniaxial heterostrain can be extended into the Hofstadter regime [42]. Here we also incorporate biaxial heterostrain, lattice relaxation, and electron-hole asymmetry so that the locations of the three van Hove points near $B=0$ on the hole side of the charge neutrality point differ from those on the electron side [43, 44], as is seen in the experiment [36, 37]. We fully parameterize the $2\times 2$ moiré deformation matrix as a function of twist angle, uniaxial heterostrain magnitude and direction, and biaxial heterostrain magnitude. We then search for structural parameters that yield bands with van Hove points at densities matching those identified in Ref. [37]. We find a unique best fit—$0.24\pm 0.02$% uniaxial heterostrain at $45\pm 4$° and $0.3\pm 0.1$% biaxial heterostrain—that places all six computed van Hove densities within our experimental bounds. Nearby moiré deformation matrices yield similar high-field Hofstadter spectra, so they do not further constrain our determination of structural parameters.
@@
Our model suggests that even without electron-electron interactions, gapped ground states with $t=0$ may exist in which the net spin of all filled bands is non-zero. We identify two such regimes, marked with B and C respectively in Fig. 3A. In each case, just below half flux the Zeeman splitting exceeds the bandwidth of a subband separating two gaps associated with Středa lines that cross at half flux. The Středa parameters for the crossing gaps in these two regimes are (B) (-4, 4) and (0, -4) (C) (0, 4) and (4, -4). The configuration at each of (B) and (C) is a quantum spin Hall insulator, though with not one but two pairs of counter-propagating modes. We might therefore expect quantized longitudinal resistance $R_{xx}=h/{4e^{2}}$ and zero Hall resistance. The Hall resistance is indeed zero below half flux near density $n/n_{s}=\mp 2$ (Fig. 3B and C lower panel, though the zero plateau in C is offset to slightly higher density). At $n/n_{s}=\mp 2$, in longitudinal resistance a narrow feature emanates downward from half flux(Fig. 3B and C upper panel). However, the longitudinal resistivity is not quantized as expected from our model, but is instead very low, often below our measurement floor of roughly 1 $\mathrm{\Omega}$ (in other contact pairs we see similar features but the minima are higher, see Sec. S9). We see similar vertical features in $R_{xx}$ at quarter flux where the ($\pm$8, $\mp$8) and (0, $\mp$8) gaps intersect (see Fig. S4). Given this unexpected behavior of longitudinal resistivity we cannot prove that some or all of these states in our experiment are QSH-like. If they indeed are QSH-like, our measurements would suggest that back-scattering is strongly suppressed not only along an edge (as expected in an ideal QSH system) but also at the ohmic contacts that serve as voltage probes, either because the edge modes do not enter these contacts or because the efficient spin relaxation normally expected within such contacts is somehow blocked.
@@ IV Landau level reset at half flux
This reset behavior is not the same bending behavior noted in Ref. [36] and cannot be understood in a non-interacting model of rigid bands. It is reminiscent of the spontaneous flavor polarization at integer filling in magic-angle twisted graphene structures at zero field [2, 3, 47, 48]. However, though spontaneous polarization into an isospin flavor can lead to reset behavior, it does not explain the observation that a dip in longitudinal resistance coincident with quantized Hall resistance follows a nonlinear trajectory in the space of filling and flux; an incompressible Chern gap must follow the appropriate Středa relation.
Figure 4: Landau level reset at half flux. (A) Zoomed segment of Fig. 1 showing the Landau levels bending and resetting at half flux. The numbers ±$\pm 6$ and ±$\pm 10$ are the slopes ($t$ values) of the faint lines $(3,\pm 6)_{1/2}$ and $(3,\pm 10)_{1/2}$ to their left. (B) Inverse Hall resistance showing that quantization of the Hall resistance tracks the bending dips in longitudinal resistance. Detail of all contact pairs in Figs. S10, S11, and S12.
For each of the four combinations of spin and valley, the two zero-field moiré minibands split into four magnetosubbands at half flux, reflecting a unit cell doubling in the Hofstadter spectrum. Because of this unit cell doubling, integer filling of the highest-energy magnetosubband (two valley-degenerate subbands split by Zeeman) corresponds to fillings $n/n_{s}=2.5$, $3$, $3.5$, and $4$ referenced to the original unit cell. At $n/n_{s}=3$, if the carriers within this band fully polarize into half of the available flavors, we expect a total Chern number of all filled bands of 2. One can see this from the fact that this gap must be the same as the (2, 2) Hofstadter subband ferromagnet state emanating from zero field [9, 24]. Landau levels emanating from half flux carry $C=2$, therefore we would expect a four-fold degenerate fan with an offset of 2 ($t=\pm 2,6,10,\ldots$) emanating from $n/n_{s}=3$. In our measurements, the Landau levels emanating from the reset point are indeed fourfold degenerate, with $(3,\pm 6)_{1/2}$ and $(3,\pm 10)_{1/2}$ being the only visible gaps. Here, we use the subscript $1/2$ to indicate we are using $s$ to reference a density offset at half of a flux quantum per moiré unit cell.

### v3

@@ Abstract
Fractional Chern insulators (FCIs)—the lattice analog of fractional quantum Hall states—form as fractionalized quasiparticles emerge in a partially-filled Chern band. This fractionalization is driven by the interplay of electronic interaction and quantum geometry of the underlying wavefunctions. Bilayer graphene with an interlayer twist near the magic angle of 1.1°  hosts diverse correlated electronic states at zero magnetic field. When the twist angle exceeds 1.3°, the electronic bandwidth is sufficient to suppress the zero-field correlated states. Yet applying a magnetic field can restore the importance of electron-electron interactions. Here, we report strongly-correlated phases when a 1.37° twisted bilayer graphene sample is tuned to near half a magnetic flux quantum per moiré cell, deep into the Hofstadter regime. Most notably, well-quantized odd-denominator FCI states appear in multiple Hofstadter subbands over unusually large ranges of density. We also observe a bending and resetting of the Landau minifan reminiscent of behavior commonly seen in magic-angle samples near integer filling at low magnetic field.
@@ Introduction
Twisted bilayer graphene (TBG), the archetypal strongly-correlated moiré heterostructure, displays astonishingly diverse correlated electronic and topological phases [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. The electronic moiré miniband structure depends sensitively and predictably on twist angle. Near the magic angle of 1.1°, the large moiré length scale enables tuning the density of charge carriers through entire minibands and the density of magnetic flux to on order one flux quantum $\Phi_{0}=e/h$ per moiré cell, where $e$ is the elementary charge and $h$ is Planck’s constant. In this “Hofstadter’s butterfly” regime [13, 14, 15, 16, 17, 18, 19, 11, 20, 21, 22, 23, 24], the energy spectrum exhibits a fractal structure with field-induced topological subbands. Gaps in the fractal spectrum appear at fluxes and densities described by Diophantine equations of the form $n/n_{s}=s+t\Phi/\Phi_{0}$ [25]. Here, $n_{s}=1/A$, $\Phi=BA$, $A$ is the moiré unit cell area, $B$ the magnetic field normal to the plane of the samples, $s$ the density offset at zero flux, and $t$ the Chern number associated with the gap. We notate these Středa lines as $(s,\ t)$. Within such a gap, the Hall conductance is expected to be quantized to $\sigma_{xy}=te^{2}/h$ [26].
Integer $t$ Středa lines have been observed in TBG at a range of twist angles. Some of these states are not describable by single-particle models, as symmetries are broken by electronic interactions [27, 7, 12, 22, 24]. Středa lines with fractional $t$ and nonzero $s$ are more exotic and are referred to as fractional Chern insulators (FCIs) [28, 29, 30, 31, 32, 33, 34, 35]. Analogous to fractional quantum Hall (FQH) states, where $s=0$, FCIs host fractionalized excitations [36]. Only a few examples of moiré materials hosting such states have been experimentally reported, in twisted MoTe2  [37, 38, 39] and multilayer twisted graphene aligned with hexagonal boron nitride (hBN)  [19, 8, 40, 41].
@@
In Refs. [44] and [45], we presented magnetotransport measurements and single-particle calculations for a TBG sample twisted to near 1.37°, where the miniband width is large enough (near 100 meV) to suppress correlated electronic states near zero magnetic field [1, 3, 44]. In this work, we further explore that same sample, now at extremely high magnetic fields, demonstrating many $B$-induced strongly-correlated states. Most notably, at magnetic flux near half a magnetic flux quantum per moiré unit cell we observe plateaus in Hall resistance corresponding to ±$\pm 8/3$ and $-8/5$, quantized to within a few tenths of a percent, along with possible $-4/3$ plateaus. The ±$\pm 8/3$ plateaus in particular extend over a strikingly broad density range, up to half an electron per moiré unit cell, rather than following a narrow Středa line. Integer quantized states have been known to extend over a broad density range, outcompeting fractional states that might be expected at those densities. This “re-entrant quantum Hall effect” has been explained as a consequence of a Wigner crystal coexisting with a filled Landau level [46, 47], a scenario originally termed a partial Hall crystal [48]. In our present work, in contrast with prior reports, the integer states are less robust and broad than the $8/3$ fractional states. In Sec. V we discuss several mechanisms that might be responsible for the novel and surprising extended FCI regions.
@@ I Longitudinal transport
In our sample, longitudinal resistivity as a function of carrier density and magnetic field between 14 and 28 T ranges from below our measurement floor of a few ohms within quantum Hall gaps to hundreds of kilohms at band edges and near charge neutrality (Fig. 1A). Here, we focus on specific contact pairs within this 20-terminal Hall bar that demonstrated the sharpest low-field transport in Ref. [44]. For our twist angle of 1.37°, $\Phi/\Phi_{0}=1$ occurs at nearly 45 T, so the field range in our high-field measurements corresponds to flux ratios between 0.31 and 0.62. We indicate the most prominent Středa lines, $(s,\ t)=(0,\,\pm 4)$, which persist without closing down to near-zero field (Fig. S2).
Figure 2: Fractional Chern insulating Hall plateaus. (A) Longitudinal resistivity from Fig. 1 replotted in gray-scale. Colored regions superimposed on the grey-scale plot mark where the Hall resistance for a neighboring contact pair is quantized to within 1% of $-8/3$ (blue) or $+8/3$ (red). A clustering algorithm has been applied to remove spurious points (see Sec. S8). The abrupt loss of quantization at $\Phi/\Phi_{0}$ between $1/3$ and $2/5$ corresponds to a pause in the measurement of several hours. This, along with the noise above $1/2$, are described in Sec. S1. (B, C) Hall resistance and longitudinal resistivity as a function of density at 18.5 T ($\Phi/\Phi_{0}=0.41$, black dashed lines in A).
For comparison, we show the computed Wannier plot in the same field range (Fig. 1B). The computed density of states $dn/d\mu$ should show patterns similar to those seen in resistivity, but we do not directly calculate transport from our model as we did in [45] for lower magnetic fields. Our model, which incorporates uniaxial heterostrain, Zeeman effect, and electron-hole asymmetry, broadly matches our experimental observations of splitting and bending behavior of Landau levels emanating from half flux and $n/n_{s}=0$. The model predicts a number of prominent gaps that we discuss in Sec. III together with associated experimental measurements.
In the Supplement we describe other striking phenomena such as Brown-Zak oscillations, symmetry-broken Chern insulating Středa lines with $(s,t)=(\pm 1/2,\pm 3$), and correlated Hofstadter ferromagnets.
@@ II Fractional Chern Insulators
In Fig. 2A, we overlay in blue and red the regions where the Hall resistance is quantized to within 1% of $-3h/8e^{2}$ and $+3h/8e^{2}$ respectively. We apply an intentionally-conservative clustering threshold algorithm so that almost all of the pockets marked with color represent well-quantized plateaus of ±$\pm 8/3$ rather than regions where the Hall resistance incidentally passes through quantized values. These fractional Hall plateaus coincide with low longitudinal resistivity (shown in gray). See Sec. S8 for details and Fig. S16 for unfiltered data.
@@
Fig. 2 panels B and C show line cuts of $R_{xy}$ and $\rho_{xx}$ at 18.5 T. The aforementioned ±$\pm 8/3$ plateaus are most prominent, but integer plateaus are visible, along with a small $-8/5$ plateau. For a discussion of the degree of quantization, other fractions, and the behavior in other contact pairs, see Secs. S9,  S10, and  S5, respectively.
@@
Figure 3: Computed Hofstadter spectrum and candidate quantum spin Hall state. (A) Computed energy levels for K valley of strained TBG ($q\leq 72$), as described in the text. Parameters: $\theta=1.35^{\circ}$, $0.24\%$ uniaxial heterostrain at 45° and $0.3\%$ biaxial heterostrain. Spin up electrons are shown in blue, and spin down electrons are shown in red. The gaps at $(s,t)=(0,-4)$ and $(4,-4)$ are labeled. $\alpha$, $\beta$, and $\gamma$ denote the Hofstadter subbands which host the largest pockets of ±$\pm 8/3$ quantization shown in Fig. 2. Two gaps $(s,t)=(0,\mp 2)$, marked B and C, are reminiscent of quantum spin Hall states. See Fig. S21 for the full range of fields. (B-C) Dependence of the longitudinal resistivity (top panels) and Hall resistance (bottom panels) as a function of density and field centered on the two quantum spin Hall candidate gaps predicted from (A). In (B), below half flux we observe a strong suppression of longitudinal resistivity at fixed density $n/n_{s}=-2$ over a range of flux, coincident with a small plateau of zero Hall resistance. In (C), we observe low longitudinal resistivity along a vertical line, however the Hall resistance is not zero at the same densities.
@@
We find that regions where the Hall resistance is (fractionally) quantized typically coincide with suppressed longitudinal resistivity in the four adjacent contact pairs (see Fig. 1 and Sec. S5). Though measurements from neighboring longitudinal probes vary subtly, neighboring Hall probes exhibit strikingly different behavior: those just 3 microns away from the main Hall pair we focus on in this manuscript exhibit poor quantization, even for integer quantum Hall states (Fig. S13), possibly due to mixing of the longitudinal resistivity into the Hall measurement because of spatial variation in twist angle  [^1] The standard approach to remove the effect of mixing is to (anti)symmetrize the longitudinal resistivity (Hall resistance) as a function of magnetic field. Antisymmetrization could help reveal well-quantized plateaus in other Hall pairs. Given the time constraints of our measurement run at the National High Magnetic Field Lab, we were unable to acquire data for opposite field polarity.. We therefore suspect that the main Hall pair of this manuscript contacts a region of unusually high moiré uniformity.
@@ III Hofstadter Model
To understand the Hofstadter bands we would expect without interactions we perform bandstructure computations within a single-particle model. Our effective continuum model from Ref. [45], which accounted for uniaxial heterostrain, can be extended into the Hofstadter regime [50]. Here we also incorporate biaxial heterostrain, lattice relaxation, and electron-hole asymmetry so that the locations of the three van Hove points near $B=0$ on the hole side of the charge neutrality point differ from those on the electron side [51, 52, 53], as is seen in the experiment [44, 45]. We fully parameterize the $2\times 2$ moiré deformation matrix as a function of twist angle, uniaxial heterostrain magnitude and direction, and biaxial heterostrain magnitude. We then search for structural parameters that yield bands with van Hove points at densities matching those identified at zero field in Ref. [45]. We find a unique best fit—$0.24\pm 0.02$% uniaxial heterostrain at $45\pm 4$° and $0.3\pm 0.1$% biaxial heterostrain—that places all six computed van Hove densities within our experimental bounds. See Sec. S12 for details.
@@
Our model suggests that even without electron-electron interactions, gapped ground states with $t=0$ may exist in which the net spin of all filled bands is non-zero. We identify two such regimes, marked with B and C respectively in Fig. 3A. In each case, just below half flux the Zeeman splitting exceeds the bandwidth of a subband separating two gaps associated with Středa lines that cross at half flux. The Středa lines that cross at (B) are parameterized by (-4, 4) and (0, -4), and those at (C) by (0, 4) and (4, -4). The configuration at each of (B) and (C) is a quantum spin Hall insulator, with not one but two pairs of counter-propagating modes. We might therefore expect quantized longitudinal resistance $R_{xx}=h/{4e^{2}}$ and zero Hall resistance. The Hall resistance is indeed zero below half flux near density $n/n_{s}=-2$ (Fig. 3B, lower), though not at $n/n_{s}=2$ (Fig. 3C, lower). At $n/n_{s}=\pm 2$, in longitudinal resistivity a narrow feature emanates downward from half flux(Fig. 3B and C, upper). However, the longitudinal resistivity is not quantized as expected from our model, but is instead very low, often below our measurement floor of roughly 1 $\mathrm{\Omega}$ (in other contact pairs we see similar features but the minima are higher, see Fig. S4). We see similar vertical features in $\rho_{xx}$ at quarter flux where the ($\pm$8, $\mp$8) and (0, $\mp$8) gaps intersect (see Fig. S2). Given this unexpected behavior of longitudinal resistivity we cannot prove that some or all of these states in our experiment are QSH-like. If they indeed are QSH-like, our measurements suggest that the edge modes do not enter the ohmic contacts that serve as voltage probes, because of the narrowness of the arms connecting to the contacts and/or the multiple densities and thus fillings in series en route to the contacts. In this scenario, the voltage read out could be an average of the electrochemical potential of ballistically-propagating left-going and right-going modes, and thus could be almost identical at successive contacts.
@@ IV Landau level reset at half flux
This reset behavior is not the same bending behavior noted in Ref. [44] and cannot be understood in a non-interacting model of rigid bands. It is reminiscent of the Landau fan resets at integer filling in magic-angle twisted graphene structures [2, 3, 56, 57, 58, 56, 22, 59, 7, 8, 9]. One possible explanation of this reset behavior in magic-angle samples is spontaneous isospin polarization. However, isospin polarization is inconsistent with our observation of a dip in longitudinal resistivity coincident with quantized Hall resistance following a nonlinear trajectory in the space of filling and flux; an incompressible Chern gap must follow the appropriate Středa relation.
Figure 4: Landau level reset at half flux. (A) Zoomed segment of Fig. 1 showing the Landau levels bending and resetting at half flux. The numbers ±$\pm 6$ and ±$\pm 10$ are the slopes ($t$ values) of the faint lines $(3,\pm 6)_{1/2}$ and $(3,\pm 10)_{1/2}$ to their left. (B) Inverse Hall resistance showing that quantization of the Hall resistance tracks the bending dips in longitudinal resistivity. Detail of all contact pairs in Figs. S8, S9, and S10.
For each of the four combinations of spin and valley, the two zero-field moiré minibands split into four magnetosubbands at half flux, reflecting a unit cell doubling in the Hofstadter spectrum. Because of this unit cell doubling, integer filling of the highest-energy magnetosubband (two valley-degenerate subbands split by Zeeman) corresponds to fillings $n/n_{s}=2.5$, $3$, $3.5$, and $4$ referenced to the original unit cell. At $n/n_{s}=3$, if the carriers within this band fully polarize into half of the available flavors, we expect a total Chern number of all filled bands of 2. One can see this from the fact that this gap must be the same as the (2, 2) Hofstadter subband ferromagnet state emanating from zero field [9, 24]. Landau levels emanating from half flux carry an extra two-fold degeneracy (see Sec. S7), therefore we would expect a four-fold degenerate fan with an offset of 2 ($t=\pm 2,6,10,\ldots$) emanating from $n/n_{s}=3$. In our measurements, the Landau levels emanating from the reset point are indeed fourfold degenerate, with $(3,\pm 6)_{1/2}$ and $(3,\pm 10)_{1/2}$ being the only visible gaps. Here, we use the subscript $1/2$ to indicate we are using $s$ to reference a density offset at half of a flux quantum per moiré unit cell.